OpenStudy (anonymous):

x+2y is greater than or equal to 8

OpenStudy (anonymous):

what are you asking?

OpenStudy (anonymous):

ii have math homework to do and we have to graph in-equations and then shade it

OpenStudy (campbell_st):

ok... so you are graphing \[x + 2y \ge 8\] find the intercepts these are the points where the line cuts the x axis and y axis do you know how to do that..?

OpenStudy (anonymous):

Kind of So Far ii HAve My X As 2,6,4,0 and my Y as 3,1,2,4

OpenStudy (campbell_st):

I guess no... just look at the line x + 2y = 8 for the x- intercept, let y = 0 so you need to solve x = 8 for the y intercepts let x = 0 so 2y = 8 or y = 4 so your line goes through the points (8, 0) and (0, 4) since the inequality is greater than or equal to the line will be solid... can you draw the line..?

OpenStudy (anonymous):

Yes i CAn

OpenStudy (anonymous):

Do You Know About Test Points?

OpenStudy (campbell_st):

can you use the draw tool to show me

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

i don't really know how to use the tool but i know my 4 is my x and 0 is the y and my line should be solid

OpenStudy (anonymous):

you've been a great help so far now can you please explain the test points

OpenStudy (campbell_st):

here is the line using the intercepts  so to test the region of the inequality pick a point say (1, 1) x = 1 and y = 1 is \[1 + 2\times 1 \ge 8\] false so (1, 1) isn't in the region..

OpenStudy (campbell_st):

which means  hope it helps